Q1 = 7.5 uC, Q2= 7.2 uc , Q3= -12.9uC ,

The component of the resultant electric force on Q3 on the X-axis is


1
Expert's answer
2021-04-19T17:44:40-0400

To be given in question

Q1=7.5"\mu c"

Q2=7.2"\mu c"

Q3=-12.9"\mu c"

To be asked in question

Electric force on x-axis

We know that

"F_{13}=k\frac {q_{1}q_{3}}{r^2}" "\rightarrow(1)"

"F_{23}=k\frac {q_{2}q_{3}}{r^2} \rightarrow(2)"

"F=\sqrt{F_{13}^2+F_{23}^2+2F_{13}F_{23}cos\theta}"

"\theta=90°"

Put "\theta" value

"F=\sqrt{F_{13}^2+F_{23}^2}\rightarrow(3)"

equation (1)put values

"F_{13}=-\frac{9\times10^9\times7.5\times10^{-6}\times12.9\times10^{-6}}{r_{13}^2}\rightarrow(4)"

"F_{13}=-\frac{0.87075}{r_{13}^2}\rightarrow(4)"

"F_{23}=-\frac{9\times10^9\times7.2\times10^{-6}\times12.9\times10^{-6}}{r_{23}^2}"

"F_{23}=\frac{-0.83592}{r_{23}^2}\rightarrow(5)"

equation (3); equation (4)and eqution (5) we can written as

"F=\sqrt(-\frac{0.87075}{r_{13}^2})^2+(\frac{-0.83592}{r_{23}^2})^2\rightarrow(6)"


equation (6)put value r13and r23

and find the resultant force